# Supervised-Contrastive Loss Learns Orthogonal Frames and Batching Matters

### v1

@@ 6.1 More related work
The simple concept behind SCL, introduced as an extension of the contrastive loss [[, e.g.,]]chen2020simple, tian2019contrastive to the fully supervised setting by [Kho+20, ], involves pulling together the normalized features of examples belonging to the same class while pushing them away from examples of other classes. Despite its simplicity, SCL offers a generalization of existing loss functions like the triplet loss [WBS05] and the N-pair loss [Soh16], and surpasses the performance of the CE loss on standard vision classification datasets [Kho+20], while also being more robust to natural corruptions in the data and less sensitive to hyperparameters. Further studies have also emphasized the observed resilience of SCL and variants alike to distribution shifts, particularly in scenarios involving training with imbalanced data [Gun+20, Jit+22, SC21, Li+22, Gun+20, Kan+21, Li+22]. These studies commonly involve training a combination of CE and SCL-type losses.
@@
In recent years, there has been a notable increase in interest regarding the properties of the embedding space and weight vectors learned through unsupervised and supervised contrastive losses, as well as CE. Of particular importance is the endeavor to uncover the fundamental distinctions between these losses, which could be essential in order to effectively harness the respective strengths of each method and formalize principled ways to combine them. Specifically, an increasing series of recent works have focused on uncovering the embedding geometry of CE, Mean-Squared Error (MSE) losses, and their variants [PHD20, Zhu+21, Zho+22, Zho+22a, LS20, Thr+22, Yar+22, Beh+23, Gao+23, Zho+22a, Yar+22, Gra+21, TB22, HPD21, Fan+21, SML23]. In addition to the scientific curiosity surrounding such studies, this exploration has paved the way for novel CE-based approaches to training DNNs on imbalanced data [Beh+23, LD23, Xie+23, Sha+23, Dan+23].
However, less attention has been devoted in the existing literature to a comparable set of comprehensive results for the Supervised Contrastive Learning (SCL) loss. [Gra+21, ] is the first study that explicitly delves into understanding the embedding geometry of SCL and comparing it to that of CE. While this study is pioneering, we note two limitations. Firstly, their investigation primarily focuses on balanced data. Secondly, to reconcile the theoretically predicted simplex ETF geometry with the observed experimental outcomes, they adopt a heuristic strategy involving shifting the center of the embeddings. However, this approach lacks theoretical support and does not result in predicting the orthogonal angle between embeddings, as depicted in Fig. 1. To tackle the first point, and inspired by suggestions in [Gra+21] that balanced data is crucial for obtaining symmetric embeddings geometries, [Zhu+22, ] have very recently proposed and investigated a clever modification of the SCL that significantly boosts performance under class imbalances. Our work addresses the two limitations identified above in the study of [Gra+21] and complements the findings of [Zhu+22] by showing that vanilla SCL can actually learn symmetric embedding structure even in the presence of imbalances.
6.1.1 Why is batch analysis “easier” for $\text{UFM}_{+}$ rather than UFM?
6.2 Detailed comparison between UFM and ${}_{+}$
Within this section, we provide an in-depth analysis, expanding upon our earlier discussions, of the disparities between the predictions generated by the UFM, commonly utilized in previous studies, and the refined version $\text{UFM}_{+}$ that we employ in this work to study the embedding geometry of SCL.
6.2.1 Centering heuristic
6.2.2 Centered OF is simplex ETF
Definition 5 (Simplex ETF).
[…]
Figure 9: Comparing the convergence of embeddings geometries for SCL (in green) and CE (in blue) on balanced CIFAR10 with ResNet-18. Solid lines track the distance of uncentered embeddings to OF and dashed lines track the distance of centered embeddings to simplex ETF. For CE, the learning rate and weight decay and batch size are set according to [PHD20]. Note the convergence of SCL (solid green) to its implicit geometry is noticeably better compared to CE (dashed blue).
In this paper, we show instead that the global optimizer of $\text{UFM}_{+}$ is an OF. Remarkably, this result holds true regardless of the label-imbalance profile. However, for the purpose of comparison with the findings of [Gra+21], let’s specifically consider the scenario of balanced data. In this case, an intriguing question arises:
Specifically for balanced data, how does our discovery that embeddings converge to an OF compare to the previous finding by [Gra+21] that embeddings converge to an ETF?
The answer to this question lies on the centering trick discussed in Section 6.2.1. Specifically, stating that embeddings form an OF implies that centered embeddings form an ETF. This straightforward fact is formalized in the following lemma.
Lemma 6.1.
Lemma 6.1 states that based on our finding that embeddings form an OF, we can deduce that centered embeddings form an ETF. This conclusion aligns with the analysis conducted by [Gra+21] on the UFM. However, the UFM analysis itself does not provide information regarding the necessity of the centering technique, which remains heuristic. Additionally, it is important to note that it is unclear how to establish that (uncentered) embeddings form an OF if we initially know that centered embeddings form an ETF. This is due to the unknown vector used for centering, which cannot be determined. This issue of identifiability is the underlying reason why the angle prediction eventually made by [Gra+21] in Fig. 1 deviates from the true value of ninety degrees, which corresponds to the OF.
6.2.3 UFM can fail to predict the true geometry
In the previous section, we demonstrated that for balanced data, the solution found by the UFM does not predict the true geometry of the embeddings (i.e., OF), but it can predict the geometry of the embeddings after the application of a global-mean centering heuristic (i.e., ETF). Naturally this raises the following question:
[…]
Lemma 6.2.
[…]
Remark 4.
While [Zhu+22, ] acknowledge that the global optima of the UFM do not conform to the ETF geometry in the presence of class imbalance, they do not provide a formal proof to support this claim. In this context, Lemma 6.2 fulfills this purpose.
6.3 Comparing implicit Geometries: SCL vs CE – A summary
  1. The SCL geometry is robust to label-imbalances. Specifically, the geometry of SCL embeddings exhibits symmetry consistently, whereas imbalances introduce distortions in the implicit geometry of CE embeddings [Fan+21, Thr+22].
@@ 7 Concluding remarks
We have shown consistent empirical and theoretical evidence that SCL learns training embeddings that converge to the OF geometry irrespective of the level of class imbalance. We believe this finding contributes a unique result to the growing literature of neural-collapse / implicit-geometry phenomena. For balanced data, our results demystify the heuristic centering of the theoretical ETF geometry, as implemented by [Gra+21, ] in their attempt to match experiments where embeddings are non-negative. For imbalanced data, our results stand as the first explicit characterization of the geometry, dispelling suspicions that imbalances alter the geometry [Zhu+22]. In both cases, these advancements are achieved by analyzing a refined UFM model, $\text{UFM}_{+}$, that closely aligns with experimental conditions by constraining embeddings to the non-negative orthant. The refined model also leads to new findings about the intricate role of batching in SCL optimization, that may be of independent interest. However, like many studies on neural-collapse phenomena, our results also have a common limitation. They only provide insights into the behavior of deep neural networks during training. On the other hand, there is a related line of recent research that explores algorithmic modifications to SCL [Gun+20, Jit+22, SC21, Li+22, Gun+20, Kan+21, Li+22], often rooted but not explicitly connected to geometric principles alike, to enhance generalization under imbalances. Ultimately, we envision the two research streams merging through the ongoing exchange of ideas.
@@
Remark 5 (Comparison to analysis of [Gra+21, ]).
[…]
Consider a simple example: $k=d=3,y_{1}=1,y_{2}=2,y_{3}=3,B_{1}=\{1,2\},B_{2}=\{1,3\},B_{3}=\{2,3\},\mathcal{B}=\{B_{1},B_{2},B_{3}\}$. Let us identify the optimal embeddings for the two scenarios with and without non-negativity constraints on the embedding coordinates.
  1. 1.
    Without non-negativity (UFM): the optimal embedding configurations can be shown to be ETF with $2$ vectors for every batch, which is simply an antipodal structure. In other words, the batch-wise optimal solutions are $\mathbf{h}_{1}=-\mathbf{h}_{2},\mathbf{h}_{1}=-\mathbf{h}_{3}$ and $\mathbf{h}_{2}=-\mathbf{h}_{3}$, for the batches $B_{1},B_{2}$ and $B_{3}$, respectively. However, the batch-wise optimal solutions are infeasible due to their contradictory nature. From [Gra+21], it can be deduced that the overall optimal configuration is instead an ETF with $3$ vectors. This example underscores the difficulty in optimizing the loss over every batch separately in case of SCL without non-negativity.
  2. 2.
[…]
@@
Appendix C Proofs for Section 6.2
C.1 Proof of Lemma 6.1
@@
Appendix D Additional experimental results and discussion
D.1 Details on the main experimental setup
@@
In our deep-net experiments, we focus on two common network architectures, ResNet and DenseNet. Specifically, we use ResNet-18 and DenseNet-40 with approximately $11$ million and $200$ thousand trainable parameters, respectively. For both models, we replace the last layer linear classifier with a normalization layer (normalizing such that $\|\mathbf{h}_{i}\|=1$ for $i\in[n]$) of feature dimension $d=512$. Specifically, following [Gra+21], we directly optimize the normalized features that are then used for inference. (We note this is slightly different compared to [Kho+20], where the authors train the output features of a projection head, then discard this projection head at inference time). We remark that both ResNet and DenseNet architectures include ReLU activations before the final output, which enforces a non-negativity on the learned embeddings. Resnet-18 models with CIFAR10, MNIST and FashionMNIST have been trained for 350 epochs with a constant learning rate of $0.1$, no weight decay, batch size of $1024$, and SCL temperature parameter $\tau=0.1$ (consistent with the choice of $\tau$ made in [Kho+20, Gra+21]). For all experiments with other models or datasets, we provide experimental details in the following sections. All models, regardless of architecture or dataset, have been trained on a single Tesla V100 GPU machine.
D.2 Additional geometric analysis
@@ D.2.1 Neural Collapse
[…]
Figure 10: Neural Collapse metric $\beta_{\text{NC}}:=\operatorname{tr}(\Sigma_{W}\Sigma_{B}^{\dagger})/k$ for the corresponding ResNet-18 experiments in Fig. 2. Values are usually on the order of $10^{-3}$ suggesting strong convergence of embeddings to their class means.
D.2.2 Angular convergence
@@
D.2.3 Embedding heatmaps
As a qualitative measure, we have generated heatmaps that visually represent the learned embedding geometries; see Figs. 3, 12, 13. Specifically, we generate heatmaps of the Gram-matrices $\mathbf{G}_{\mathbf{M}}=\mathbf{M}^{\top}\mathbf{M}$ and $\mathbf{G}_{\mathbf{H}}={\mathbf{H}}^{\top}{\mathbf{H}}$. In Fig. 3 we train ResNet-18 with the full MNIST dataset. In Fig. 12 we run on a subset of the dataset ($n=10000$ total examples) with a batch size of 1000. Specifically for Fig. 12, when optimizing with CE loss we modify the network (ResNet-18) such that it has a normalization layer before the linear classifier head. We consider this additional setting to allow for a comparison where CE features are constrained to the unit sphere akin to our SCL experiments. Lastly, in Fig. 13 we plot the learned features Gram-matrix $\mathbf{G}_{\mathbf{H}}$ for a ResNet-18 trained on CIFAR10 ($n=10000$ total examples) with a batch size of $1000$. [^6] In both Fig. 13 and Fig. 12 no batch duplication is used as described in Sec. 3. This heatmap qualitatively shows a more complete picture as we are plotting $\mathbf{G}_{\mathbf{H}}={\mathbf{H}}^{\top}{\mathbf{H}}$ rather than $\mathbf{G}_{\mathbf{M}}$, thus simultaneously illustrating both Neural Collapse and convergence to the k-OF structure.
D.2.4 Experiments with MLPs
In Fig. 16 we run experiments with a simple 6 layer multilayer perceptron (MLP) to further explore the impact of model complexity on geometric convergence. The MLP includes batch normalization and ReLU activation between each layer. Each layer has 512 hidden units. We train the model with a batch size of 1000 with random reshuffling at each epoch. Furthermore, we train under $R$-STEP imbalanced MNIST. No batch duplication was used. All other aspects of the implementation are as described in Sec. D.1. As shown in Fig. 16 all metrics $\Delta_{\mathbf{G}_{\mathbf{M}}}$, $\beta_{\text{NC}}$, and $\text{Ave}_{c\neq c^{\prime}}\text{ }\alpha_{\text{sim}(c,c^{\prime})}$ indicate strong convergence to the OF geometry, irrespective of imbalance ratio $R$.
@@ D.2.4 Experiments with MLPs
D.3 Optimization dynamics
D.3.1 Loss convergence

### v2

@@ 6 More Related Work
The simple concept behind SCL, introduced as an extension of the contrastive loss [Che+20, TKI19, e.g.,] to the fully supervised setting by [Kho+20, ], involves pulling together the normalized features of examples belonging to the same class while pushing them away from examples of other classes. Despite its simplicity, SCL offers a generalization of existing loss functions like the triplet loss [WBS05] and the N-pair loss [Soh16], and surpasses the performance of the CE loss on standard vision classification datasets, while also being more robust to natural corruptions in the data and less sensitive to hyperparameters [Kho+20].
@@
In recent years, there has been a notable interest regarding the properties of the embedding space and weights learned through unsupervised/supervised contrastive losses, as well as CE. Of particular importance is the endeavor to uncover the fundamental distinctions between these losses to effectively harness each of their strengths and formalize principled ways to combine them. Specifically, starting by [PHD20], an increasing series of recent works have focused on uncovering the embedding geometry of CE, Mean-Squared Error (MSE) losses, and their variants. We highlighted some of these works in Sec. 1 while numerous other works contribute to these investigation [Zho+22, Zho+22a, Yar+22, Gao+23, HPD21, SML23, e.g.,]. In addition to the scientific curiosity surrounding such studies, this exploration has paved the way for novel CE-based approaches to training DNNs on imbalanced data [Beh+23, LD23, Xie+23, Sha+23, Dan+23, Yan+22].
However, less attention has been devoted in the existing literature to a comparable set of results for the SCL. [Gra+21, ] is the first to study the embedding geometry of SCL for balanced data and comparing it to that of CE. To tackle the imbalances, and inspired by [Gra+21, ] that suggests balanced data is crucial for obtaining symmetric embeddings, [Zhu+22, ] have recently proposed and investigated a modification of the SCL that boosts performance under class imbalances. Our work extends and complements these two studies by showing that SCL with ReLU can actually learn symmetric embedding structure even in the presence of imbalances. While preparing the manuscript of our paper, we became aware of a recent work [Cho+23], where the authors consider the mini-batch optimization of the unsupervised contrastive loss and propose a loss-based choice of mini-batches to speed up convergence to ETF. Since their setting is different, it would be interesting to see how our ideas can benefit their findings and vice versa.
@@ 7 Concluding remarks
We have shown consistent empirical and theoretical evidence that SCL learns training embeddings that converge to the OF geometry irrespective of the level of class imbalance. We believe this finding contributes a unique result to the growing literature of neural-collapse / implicit-geometry phenomena. For balanced data, we extend the contributions of [Gra+21, ] to the case in presence of ReLU activations. Further, our results identify exact conditions on the mini-batch selection to achieve the OF geometry, allowing for a wider set of possibilities in batching than found by [Gra+21] for SCL without ReLU. For imbalanced data, our results are the first explicit characterization of the geometry, concluding that imbalances do not alter the geometry if a ReLU activation is added at the last layer. In both cases, these advancements are achieved by analyzing a refined UFM model, $\text{UFM}_{+}$, that closely aligns with experimental conditions by constraining embeddings to the non-negative orthant. This model also leads to new findings about the intricate role of batching in SCL optimization, that may be of independent interest. A future investigation of exact characterization of the geometry with imbalances and without ReLU can help complete the understanding of the implicit-geometry of SCL. Although the majority of implicit geometry characterizations in the literature require $d>k$, it is interesting to extend our findings to settings where $d<k$ following the steps of [Gao+23] for the CE loss. Likewise, our finding regarding the crucial role of batching in geometry opens the door to further investigations aimed at devising efficient batch strategies in cases with a large number of classes. Finally, like most studies on neural-collapse phenomena, our results share a common limitation: they only provide insights into the behavior during training. On the other hand, there is a line of research that explores algorithmic modifications to SCL [Gun+20, Jit+22, SC21, Li+22, Kan+21, Li+22], often rooted but not explicitly connected to geometric principles, to enhance generalization under imbalances. Ultimately, we envision the two research streams merging through the ongoing exchange of ideas.
@@
B.2 Batch analysis for $\text{UFM}_{+}$ vs UFM
Appendix C Proofs for Section D.1
C.1 Proof of Lemma D.1
@@ B.2 Batch analysis for $\text{UFM}_{+}$ vs UFM
Appendix D Additional Discussion
D.1 Detailed comparison between UFM and ${}_{+}$
Within this section, we provide an in-depth analysis of the disparities between the predictions generated by the UFM, commonly utilized in previous studies, and the refined version $\text{UFM}_{+}$ that we employ in this work to study the embedding geometry of SCL.
D.1.1 Centering heuristic
D.1.2 Centered OF is simplex ETF
Definition 5 (Simplex ETF).
[…]
Figure 7: Convergence of embeddings geometries for SCL (in green) and CE (in blue) on balanced CIFAR10 with ResNet-18. Solid lines track the distance of uncentered embeddings to OF and dashed lines track the distance of centered embeddings to simplex ETF. For CE, the learning rate and weight decay and batch size are set according to [PHD20]. Note the convergence of SCL (solid green) to its implicit geometry is noticeably better compared to CE (dashed blue).
In this paper, we show instead that the global optimizer of $\text{UFM}_{+}$ is an OF. Remarkably, this result holds true regardless of the label-imbalance profile. However, for the purpose of comparison with the findings of [Gra+21], let us specifically consider the scenario of balanced data. In this case, an intriguing question arises:
Specifically for balanced data, how does our discovery that embeddings with ReLU converge to an OF compare to the previous finding by [Gra+21] that embeddings without ReLU converge to an ETF?
The answer to this question lies on the centering trick discussed in Section D.1.1. Specifically, stating that embeddings form an OF implies that centered embeddings form an ETF. This straightforward fact is formalized in the following lemma.
Lemma D.1.
Lemma D.1 states that based on our finding that embeddings form an OF, we can deduce that centered embeddings form an ETF. This conclusion aligns with the analysis conducted by [Gra+21] on the UFM (See Fig. 7 for a representative comparison). However, the UFM analysis itself does not provide information regarding the necessity of the centering technique, which remains heuristic. Additionally, it is important to note that it is unclear how to establish that (uncentered) embeddings form an OF if we initially know that centered embeddings form an ETF. This is due to the unknown vector used for centering, which cannot be determined.
D.1.3 UFM can fail to predict the true geometry
In the previous section, we demonstrated that for balanced data, the solution found by the UFM does not predict the true geometry of the embeddings (i.e., OF) in presence of ReLU, but it can predict the geometry of the embeddings after the application of a global-mean centering heuristic (i.e., ETF). Naturally this raises the following question:
[…]
Lemma D.2.
[…]
D.2 Comparing implicit Geometries: SCL vs CE – A summary
  1. The SCL geometry with ReLU is robust to label-imbalances. Specifically, the geometry of SCL embeddings exhibits symmetry consistently, whereas imbalances introduce distortions in the implicit geometry of SCL without ReLU and also for CE [Fan+21, Thr+22].
Appendix E Additional experimental results and discussion
E.1 Details on the main experimental setup
@@
In our deep-net experiments, we focus on two common network architectures, ResNet and DenseNet. Specifically, we use ResNet-18, ResNet-34 and DenseNet-40 with approximately $63$, $11$ million and $200$ thousand trainable parameters, respectively. For all models, we replace the last layer linear classifier with a normalization layer (normalizing such that $\|\mathbf{h}_{i}\|=1$ for $i\in[n]$) of feature dimension $d=512$. Specifically, following [Gra+21], we directly optimize the normalized features that are then used for inference. (We note this is slightly different compared to [Kho+20], where the authors train the output features of a projection head, then discard this projection head at inference time). We use this projection head when studying NCC test accuracy. In particular, we add a 2 layer non-linear MLP with with and without ReLU to report the test accuracies provided in Tab. 1, Tab. 2 and , Tab. 3. Following [Kho+20], in addition to the NCC test accuracy on post-projection feature (final output), we evaluate the accuracies on the pre-projection features (without input to projector) as well. We remark that both ResNet and DenseNet architectures include ReLU activations before the final output, which enforces a non-negativity on the learned embeddings. Resnet-18 models with CIFAR10, MNIST have been trained for 350 epochs with a constant learning rate of $0.1$, no weight decay, batch size of $1024$, and SCL temperature parameter $\tau=0.1$ (consistent with the choice of $\tau$ made in [Kho+20, Gra+21]). For CIFAR100 and Tiny-ImageNet experiments, we train a ResNet-34 under a similar setup, for 500 epochs, (with batch binding in Fig. 2). For all experiments with other models or datasets, we provide experimental details in the following sections. All ResNet-18 and DenseNet models have been trained on a single Tesla V100 GPU machine and all ResNet-34 models were trained on a single machine with 2-4 Tesla V100 GPU, depending on experiment.
E.2 Details on Fig. 2 Heatmaps
Since this particular experiment provides the most insight into the geometric behaviour of features, we have provided further explanation of the experimental setup and method. ResNet-18 models were trained for 350 epochs on MNIST using CE, SCL and SCL + ReLU under difference imbalance ratios. For SCL and SCL + ReLU we extract the features, calculate the class means and plot a heatmap of the gram matrices $G_{M}$. For CE, following previous work by [Pap18, Thr+22, Zhu+21] we center the class means by the global mean, ie $\overline{\mu_{c}}=\mu_{c}-\frac{1}{K}\sum_{i=1}^{K}\mu_{i}$ and then proceed to calculate $G_{M}$. In order to have a fair comparison, we normalize all gram matrices by dividing by the matrix maximum $\overline{G_{M}}=\frac{G_{M}}{\text{max}_{i,j}|G_{\mu}^{i,j}|}$. No projection layer was implemented for these results. The heatmaps were included to provide a clear visual, emphasizing the impact of ReLU on achieving symmetric geometry under imbalance. Each cell $[i,j]$ of the $G_{M}$ matrix represents the inner product $\mu_{i}^{T}\mu_{j}$, rather a normalized value representing the angle between the two class centers. In addition, as SCL includes a normalization layer, the diagonals $G_{M}[i,i]$ would be closer to zero as the features exhibit stronger neural collapse properties.
E.3 Additional geometric analysis
[…]
Figure 8: Neural Collapse metric $\beta_{\text{NC}}:=\operatorname{tr}(\Sigma_{W}\Sigma_{B}^{\dagger})/k$ for the corresponding experiments in Fig. 2. Values are usually on the order of $10^{-2}$ suggesting strong convergence of embeddings to their class means.
@@ E.3.1 Neural Collapse
E.3.2 Angular convergence
@@
E.3.3 Embedding heatmaps
As a qualitative measure, we have generated heatmaps that visually represent the learned embedding geometries; see Figs. 1, 10, 11. Specifically, we generate heatmaps of the Gram-matrices $\mathbf{G}_{\mathbf{M}}=\mathbf{M}^{\top}\mathbf{M}$ and $\mathbf{G}_{\mathbf{H}}={\mathbf{H}}^{\top}{\mathbf{H}}$. In Fig. 1 we train ResNet-18 with the full MNIST dataset. In Fig. 10 we run on a subset of the dataset ($n=10000$ total examples) with a batch size of 1000. Specifically for Fig. 10, when optimizing with CE loss we modify the network (ResNet-18) such that it has a normalization layer before the linear classifier head. We consider this additional setting to allow for a comparison where CE features are constrained to the unit sphere akin to our SCL experiments. Lastly, in Fig. 11 we plot the learned features Gram-matrix $\mathbf{G}_{\mathbf{H}}$ for a ResNet-18 trained on CIFAR10 ($n=10000$ total examples) with a batch size of $1000$. [^7] In both Fig. 11 and Fig. 10 no batch duplication is used as described in Sec. 3. This heatmap qualitatively shows a more complete picture as we are plotting $\mathbf{G}_{\mathbf{H}}={\mathbf{H}}^{\top}{\mathbf{H}}$ rather than $\mathbf{G}_{\mathbf{M}}$, thus simultaneously illustrating both Neural Collapse and convergence to the k-OF structure.
E.3.4 Experiments with MLPs
In Fig. 14 we run experiments with a simple 6 layer multilayer perceptron (MLP) to further explore the impact of model complexity on geometric convergence. The MLP includes batch normalization and ReLU activation between each layer. Each layer has 512 hidden units. We train the model with a batch size of 1000 with random reshuffling at each epoch. Furthermore, we train under $R$-STEP imbalanced MNIST. No batch duplication was used. All other aspects of the implementation are as described in Sec. E.1. As shown in Fig. 14 all metrics $\Delta_{\mathbf{G}_{\mathbf{M}}}$, $\beta_{\text{NC}}$, and $\text{Ave}_{c\neq c^{\prime}}\text{ }\alpha_{\text{sim}(c,c^{\prime})}$ indicate strong convergence to the OF geometry, irrespective of imbalance ratio $R$.
@@ E.3.4 Experiments with MLPs
E.4 Optimization dynamics
E.4.1 Loss convergence